OpenStudy (zugzwang):

please help: −6𝑎 + 3 = −3(2𝑎 − 1)

OpenStudy (alivejeremy):

Is many solutions a choice? :D

OpenStudy (zugzwang):

I just need one that will work. As long as it works its acceptable haha(:

OpenStudy (alivejeremy):

So there aren't no choices?

OpenStudy (zugzwang):

No, sadly not :'( RIP

OpenStudy (zugzwang):

uggghhhhhhhh if I could figure it out I wouldn't of posted it, now would I? xD lol. Im doing a thing with 50+ questions and i've been stuck on about 4.

OpenStudy (mathmale):

You haven't explained the instructions for this problem: −6𝑎 + 3 = −3(2𝑎 − 1) I assume your goal is to "solve for a." If so, group all the "a" terms on the left side and all the other terms on the right side. Know what to do first?

OpenStudy (triciaal):

the first one like alivejeremy indicated will be many solutions the left side = the right side so any value of a will make the equation true

OpenStudy (mathmale):

−6𝑎 + 3 = −3(2𝑎 − 1) becomes what? after multiplication. Pls do this mult. now.

OpenStudy (zugzwang):

Kind of? I understand that's what you're supposed to do, but I fear i've been doing it incorrectly. Okay, so then it would be -6a+3=-6a -1.

OpenStudy (mathmale):

Jeremy and Tricia are ahead of me, and both have made good contributions here. I would like for you, @zugzwang, to show your work. Thank you for having done that. Please check your calculations on the right side; you've overlooked something important.

OpenStudy (mathmale):

Your -6a+3=-6a -1 is partially correct, partially wrong. Can you find the error? If so, pls fix it.

OpenStudy (zugzwang):

I canot find the error on the right side. Im confused to what I have overlooked.. Can you please explain ?

OpenStudy (zugzwang):

*cannot

OpenStudy (mathmale):

Try multiplying out "−3(2𝑎 − 1)" again. Use the distributive property of multiplication here.

OpenStudy (zugzwang):

Okay, so then: -6a - -3?

OpenStudy (mathmale):

Please don't write "- -3" ... you must do the actual multiplication. What is (-)(-1)?

OpenStudy (mathmale):

Again, multiply out "−3(2𝑎 − 1)"

OpenStudy (mathmale):

Note that because the signs of the 2 terms inside parentheses are different now, the results of mult. will also be different from one another.

OpenStudy (mathmale):

What is -3(2a)? What is -3(-1)?